Collected 30ml of sample. Sample was clear. Blank sample only 80ml due to running out of DI H2O. Sample collected at 0800
Collected 50ml of sample. Sample was brown, with sediment settling in the bottom, may be due to collector not operating for a while. Sample collected at 0830
Finally got a sample!
Lots of fog south of the GG despite yesterday’s forecast. UCSC and LML collectors were not powered last night. Surface conditions were quite dry so not sure any sample could have been collected but visibility was somewhat impaired — thick fog.
There was monsoonal flow and moisture associated with hurricane Dolores on 7/18, 7/19, and 7/20. From reading the forecast discussion, it appears that sites south of the Golden Gate will have a poorly consolidated marine layer and probably no stratus tonight 7/20. North of the GG, stratus will probably set in as the disturbance to the south is having less of an effect.
Switched to RH-only trigger > 99%.
Switched triggering to RH-only > 99%
Heavy fog last night but collector came on at 8 am until 10 am and got a few drops of sample. Triggering working properly from tipping bucket but am concerned that it triggers too late. Changing to Rh only trigger > 95%.
sample volume 100mL, water quite cloudy with a layer of grime settled on the bottom despite yesterday’s cleaning (there was lot’s of traffic last night and the collector came on well before dark), over cast when jar collected at 8:45am on 7/9/2015, rain developed later in the day. There was lots of new grime and dirt in the tray (more than previous day prior to that cleaning). Collector was cleaned, and is currently off because RH is high enough to trigger (100%), but it is not foggy. entry made at 4:10pm 7/9/2015. Will turn on at 5:00pm.
It is likely that tonight’s sample will be a mix of rain and fog. Cool! Also, I enjoyed seeing the sample of fog in the jar from my breakfast table this morning!
Peter got the active fog collector up and running very early this morning, and we were able to get our first sample today! The webcam is back online as well after an offline stint. It has been mostly fogless and dry at Pepperwood, which has above the marine layer and below the clouds for most of the early summer. This morning there was a deep marine layer with classic tiny droplets. We missed one sig fog event on Monday 7/6 due to malfunction.
sample volume 75mL, water clear, general weather descriptions, sample collected at 8:45am on 7/8/2015. There was grime and dirt in the tray, due to the long period since the last clean. Collector was cleaned and reset and program modified to trigger only when the RH reaches 99.5. entry made at 2:10pm 7/8/2015.
*A small white plastic component fell off the door (placed it in my toolbox). It looks like it was 1 of 2 components holding the wiring that runs up to the solenoid. Should be fine without it.